# SARChat-Bench-2M: A Multi-Task Vision-Language Benchmark for SAR Image Interpretation

### v2

@@ Abstract
As a powerful all-weather Earth observation tool, synthetic aperture radar (SAR) remote sensing enables critical military reconnaissance, maritime surveillance, and infrastructure monitoring. Although Vision language models (VLMs) have made remarkable progress in natural language processing and image understanding, their applications remain limited in professional domains due to insufficient domain expertise. This paper innovatively proposes the first large-scale multimodal dialogue dataset for SAR images, named SARChat-2M, which contains approximately 2 million high-quality image-text pairs, encompasses diverse scenarios with detailed target annotations. This dataset not only supports several key tasks such as visual understanding and object detection tasks, but also has unique innovative aspects: this study develop a visual-language dataset and benchmark for the SAR domain, enabling and evaluating VLMs’ capabilities in SAR image interpretation, which provides a paradigmatic framework for constructing multimodal datasets across various remote sensing vertical domains. Through experiments on 16 mainstream VLMs, the effectiveness of the dataset has been fully verified. The project will be released at https://github.com/JimmyMa99/SARChat.
@@
11footnotetext: These authors contributed equally to this work.22footnotetext: Corresponding author. Xiayang Xiao and Haipeng Wang
@@ 2.2 Remote Sensing Vision-Language Datasets
Remote sensing datasets are essential for intelligent interpretation models. Existing datasets such as UCM Captions Qu et al. (2016b), Sydney Captions Qu et al. (2016a), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), and RSVG (Zhan et al., 2023) provide preliminary resources for studying the correlation between remote sensing images and text. However, these datasets are limited not only in scale but also in modality, containing only optical images without SAR data, leaving SAR interpretation capabilities largely unexplored. Although large-scale datasets like MillionAID (Long et al., 2021), FMoW (Christie et al., 2018), and BigEarthNet (Sumbul et al., 2019) exist, they lack text-image pairs. The RS5M dataset (Zhang et al., 2023), containing 5 million image-text pairs, is still limited to optical images. The MMRS-1M dataset (Zhang et al., 2024), which covers optical, infrared, and SAR modes, has a very low proportion of SAR image-text data. Therefore, this paper constructs the SARChat-2M dataset, which focuses on SAR images and contains over 2 million image-text pairs, covering tasks such as classification, detection, caption generation, VQA, and visual grounding.
@@ 3.1.1 Dataset Overall
As shown in Figure 2, we propose SARChat-Bench-2M, a multi-task benchmark dataset for SAR images, comprising 2 million multimodal dialogue samples (1,836,912 train and 226,636 test samples) to ensure robust model training and evaluation.
The dataset covers six core SAR image analysis tasks: classification, fine-grained description, instance counting, spatial grounding, cross-modal identification, and referring. Based on the SARDet-100K dataset (Dai et al., 2024), it incorporates multimodal adaptations and enhanced language annotations from ten established SAR detection benchmarks such as AIR-SARShip(1.0&2.0)(Wang et al., 2019a), HRSID(Wei et al., 2020), MSAR(Chen et al., 2022), SADD(Zhang et al., 2021a), SAR-AIRcraf(Zhirui et al., 2023), ShipDataset(Wang et al., 2019b), SSDD(Zhang et al., 2021b), OGSOD(Wang et al., 2023), and SIVED(Lin et al., 2023). The dataset establishes image-text correspondences across six semantic categories (ships, tanks, bridges, ports, aircraft, and automobiles) through cross-modal learning, resulting in 2 million curated annotations.
This study designs the first evaluation framework for SAR remote-sensing visual-language models, addressing three key requirements: multi-task supervised pre-training, cross-domain adaptation between SAR and language domains, and comprehensive performance assessment using standardized metrics.
@@ 3.1.2 Task Definition
(4) Spatial Grounding: Spatial relationship interpretation enables comprehensive analysis of target distributions in SAR imagery, focusing on positions, distances, and directional relationships among multiple targets. This capability is particularly vital for military reconnaissance scenarios, where precise spatial understanding directly impacts operational effectiveness.
@@
(6)Referring: This is a reverse-reasoning task. The model is required to locate specific instances in SAR images based on the target semantic categories specified by users and output abstract spatial orientations, which is of great significance for achieving rapid target retrieval in human-machine collaboration.
3.2 Task-Oriented Data Generation
Based on the characteristics of the six tasks, this study designs a multimodal dialogue data generation scheme. The specific rules and implementation logic are as follows, with detailed examples provided in the Appendix .1.4:
@@ 3.2 Task-Oriented Data Generation
Our dataset adopts a unified representation scheme across all visual-language tasks to ensure consistency and interpretability. The spatial information is uniformly encoded using the bounding box format {<$x_{1}$><$y_{1}$><$x_{2}$><$y_{2}$>}, while spatial relationships are structured through a standard 3×3 grid system (consisting of top-left, top, top-right, left, middle, right, bottom-left, bottom, bottom-right regions).
@@
Fine-Grained description evaluates the model’s structured parsing of satellite imagery through comprehensive quality control. Following our dataset definitions, we filter images below 224×224 pixels and exclude targets with area ratio R < 1% (Equation 1). Targets with aspect ratios exceeding 10:1 or out-of-bounds coordinates are removed. Size descriptions are categorized using area-ratio thresholds (small: <5%, large: >30%). We construct 40 interaction templates to accommodate multi-target scenarios.
@@ 3.3 Quantitative Analysis of Datasets
As shown in Table 1, the ship category dominates both training and test sets (46.98% and 44.38% respectively), while the harbor category represents less than 2%. A significant distribution shift is observed in the aircraft category, with a 7.53% increase in the test set compared to the training set. Categories such as cars, tanks, and bridges maintain moderate and stable proportions across both sets. This imbalanced distribution presents challenges for target detection tasks while offering opportunities to investigate solutions for handling imbalanced data.
@@
Metric Training Set Test Set Diff-Rate
Mean 1.28 1.26 -2.03%
Median 1.062 1.05 -1.69%
SD 1.18 0.91 -22.82%
As shown in Table 2, the differences in central tendency between training and test sets are minimal (mean: -2.03%, median: -1.69%). The test set exhibits a 22.82% lower standard deviation, indicating a more concentrated distribution. The key morphological distribution intervals of targets are illustrated in Appendix .1.2.
The dataset exhibits three distinct morphological categories based on aspect ratio (AR): broad-bodied (AR $\leq$ 0.67), nearly square-shaped (0.67 $<$ AR $\leq$ 1.5), and tall-bodied (AR $>$ 1.5). Detailed distribution analysis can be found in Appendix .1.2.
@@
4 Evaluation Method and Settings
This section aims to introduce the evaluation methods for six tasks. These tasks evaluate the model’s core capabilities across information processing, target localization, and semantic understanding, providing multi-dimensional insights into visual-language model performance.
@@ 4.1 Evaluation Metrics
Among them, $n_{t}$ represents the count of task $t$, $a_{m,t}$ denotes the accuracy of model $m$ on task $t$, and $T$ is the set of all tasks. The detailed calculation of $a_{m,t}$ for each task can be found in Appendix .3.
@@ 4.2 Assessment Methods
(1) Instance Counting: Compare predicted and label object counts for single-class evaluation.
@@
(6) Classification: Compare predicted and ground-truth categories to assess classification accuracy.
@@ 5.1 Implementation Details
During the fine-tuning stage, the model is trained for 1 epoch with a batch size of 4 (effective batch size 32 with gradient accumulation steps of 4). This study employs LoRA training method with rank=8 and alpha=32 on all linear layers. The learning rate is initialized at 1e-4 with a 0.1 warm-up ratio. All experiments are conducted on 2 NVIDIA A100 GPUs using bfloat16 precision and gradient checkpointing.
@@ 5.2 Benchmark Evaluation
Figure 3: Evaluation examples on SARChat-Bench. VLM predictions are shown in green/red for correct/incorrect descriptions, with the ground truth in green and the predictions in red boxes. And [Human], [Bot], and [Check] denote user input, VLMs response, and standard output, respectively.
To verify the effectiveness and practicality of the SARChat-2M dataset, this study presents SARChat-Bench, a series of benchmark experiments on trained models which are based on this dataset. As shown in Table 3, sixteen mainstream visual-language models, such as Qwen2-VL(Wang et al., 2024), InternVL2.5(Chen et al., 2024), DeepSeekVL(Lu et al., 2024), Phi-3.5-vision(Abdin et al., 2024), GLM-Edge-V(GLM et al., 2024), mPLUG-Owl3(Ye et al., 2024), Yi-VL(Young et al., 2024) and LLaVA-1.5(Liu et al., 2023). In experiments of different tasks, each model exhibits diverse characteristics:
@@ 5.4 Dialogue Visualizaion
Figure 3 demonstrates representative examples across six tasks from SARChat-Bench, displaying query-response pairs with correct answers marked in green. Examples include successful classification tasks (identifying "ship" among multiple options) and accurate instance counting (detecting "4 instances" of ships in the satellite image). These results illustrate the model’s effective performance across various SAR image understanding tasks.
@@ 6 Conclusion
The research introduces SARChat-2M, a large-scale dataset comprising two million meticulously annotated image-text pairs spanning diverse SAR scenarios. Building upon this foundation, SARChat-Bench serves as a comprehensive evaluation framework that systematically assesses visual-language models performance in the SAR domain, facilitating domain-specific knowledge integration and accelerating the development of SAR-oriented visual-language models.
@@
.1 Data structure analysis
.1.1 Word Frequency Analysis of SARChat-2M
@@
.1.2 Analysis of Datasets Composition
@@
Table 4: Analysis of Aspect Ratio of Different Types of Targets
@@
.1.3 Analysis of Task based on Datasets Composition
@@
Table 5: Task type distribution in training and test sets
@@
.2 Motivation for the Dual-benchmark Setup
@@
Table 6: Performance comparison across different model sizes
Parameters 8B 4B 2B 1B
@@
.3 Task-specific Performance Scoring
@@
.4 The Analysis of Model Size

### v3

@@ Abstract
As a powerful all-weather Earth observation tool, synthetic aperture radar (SAR) remote sensing enables critical military reconnaissance, maritime surveillance, and infrastructure monitoring. Although Vision language models (VLMs) have made remarkable progress in natural language processing and image understanding, their applications remain limited in professional domains due to insufficient domain expertise. This paper innovatively proposes the first large-scale multimodal dialogue dataset for SAR images, named SARChat-2M, which contains approximately 2 million high-quality image-text pairs, encompasses diverse scenarios with detailed target annotations. This dataset not only supports several key tasks such as visual understanding and object detection tasks, but also serves as the first visual-language benchmark in the SAR domain. Through this work, we enable and evaluate VLMs’ capabilities in SAR image interpretation, providing a paradigmatic framework for constructing multimodal datasets across various remote sensing vertical domains. Through experiments on 16 mainstream VLMs, the effectiveness of the dataset has been fully verified. The project will be released at https://anonymous.4open.science/r/SARChat-D0ED/.
@@
footnotetext: These authors contributed equally to this work.**footnotetext: Corresponding author. Xiayang Xiao and Haipeng Wang
@@ 2.2 Remote Sensing Vision-Language Datasets
Remote sensing datasets are essential for models that interpret remote sensing imagery. Existing datasets such as UCM Captions Qu et al. (2016b), Sydney Captions Qu et al. (2016a), RSICD (Lu et al., 2017), RSITMD (Yuan et al., 2022), and RSVG (Zhan et al., 2023) provide preliminary resources for studying the correlation between remote sensing images and text. However, these datasets are limited not only in scale but also in modality, containing only optical images without SAR data, leaving SAR interpretation capabilities largely unexplored. Although large-scale datasets like MillionAID (Long et al., 2021), FMoW (Christie et al., 2018), and BigEarthNet (Sumbul et al., 2019) exist, they lack text-image pairs. The RS5M dataset (Zhang et al., 2023), containing 5 million image-text pairs, is still limited to optical images. The MMRS-1M dataset (Zhang et al., 2024), which covers optical, infrared, and SAR modes, has a very low proportion of SAR image-text data. Therefore, this paper constructs the SARChat-2M dataset, which focuses on SAR images and contains over 2 million image-text pairs, covering tasks such as classification, detection, caption generation, VQA, and visual grounding.
@@ 3.1.1 Dataset Overall
As shown in Figure 2, we propose SARChat-2M, a multi-task dataset for SAR images, comprising 2 million multimodal dialogue samples (1,836,912 train and 226,636 test samples) to ensure robust model training and evaluation.
Based on the SARDet-100K dataset (Dai et al., 2024), it incorporates multimodal adaptations and enhanced language annotations from ten established SAR detection benchmarks such as AIR-SARShip(1.0&2.0) (Wang et al., 2019a), HRSID (Wei et al., 2020), MSAR (Chen et al., 2022), SADD (Zhang et al., 2021a), SAR-AIRcraf (Zhirui et al., 2023), ShipDataset (Wang et al., 2019b), SSDD (Zhang et al., 2021b), OGSOD (Wang et al., 2023), and SIVED (Lin et al., 2023). The SARChat-2M covers six semantic categories (ships, tanks, bridges, ports, aircraft, and automobiles) and supports six core SAR image analysis tasks: classification, fine-grained description, instance counting, spatial grounding, cross-modal identification, and referring. These diverse tasks are designed to enhance VLMs’ capabilities in SAR image interpretation, with 2 million carefully curated annotations for cross-modal learning.
@@ 3.1.2 Task Definition
(4) Spatial Grounding: This task challenges the model to interpret and reason about complex spatial relationships between multiple targets in SAR imagery, including their relative positions, distances, and directional relationships. The key challenge lies in accurately understanding and describing diverse spatial configurations, especially in scenes with multiple interacting objects and varying spatial layouts. The model must demonstrate precise spatial reasoning abilities while handling complex multi-target scenarios and maintaining consistent performance across different scene compositions.
@@
(6) Referring: This reverse-reasoning task challenges the model to locate specific instances in SAR images from textual descriptions. The key challenge lies in bridging semantic-visual gaps while accurately determining target spatial orientations, requiring robust cross-modal reasoning capabilities across varied scene configurations.
3.1.3 Task-Oriented Data Generation
Based on the characteristics of the six tasks, this study designs a multimodal dialogue data generation scheme. The specific rules and implementation logic are as follows, with detailed templates provided in the Appendix A.4:
@@ 3.1.3 Task-Oriented Data Generation
Our dataset adopts a unified representation scheme across all visual-language tasks to ensure consistency and interpretability. The spatial information is uniformly encoded using the bounding box format {<$x_{1}$><$y_{1}$><$x_{2}$><$y_{2}$>}, where ($x_{1}$,$y_{1}$) and ($x_{2}$,$y_{2}$) denote the top-left and bottom-right coordinates respectively. Spatial relationships are structured through a standard 3×3 grid system (consisting of top-left, top, top-right, left, middle, right, bottom-left, bottom, bottom-right regions).
To explicitly specify different task requirements, we incorporate task-specific prompts: $[count]$ for Instance Counting Task, $[grounding]$ for Spatial Grounding Task, $[identify]$ for Cross-Modal Identification, and $[refer]$ for Referring Task. These prompts help guide the model’s attention to the relevant aspects of each task.
@@
Fine-Grained description evaluates the model’s structured parsing of satellite imagery through comprehensive quality control. Following our dataset definitions, we filter images below 224×224 pixels and exclude targets with area ratio $R$ < 1% (Equation 1). Targets with aspect ratios exceeding 10:1 or out-of-bounds coordinates are removed. Size descriptions are categorized using area-ratio thresholds (small: <5%, large: >30%). We construct 40 interaction templates to accommodate multi-target scenarios. The calculation of $R$ is formulated as follows:
@@ 3.2 Quantitative Analysis of Datasets
As shown in Table 1, the ship category dominates both training and test sets (46.98% and 44.38% respectively), while the harbor category represents less than 2%. A significant distribution shift is observed in the aircraft category, with a 7.53% increase in the test set compared to the training set. Categories such as cars, tanks, and bridges maintain moderate and stable proportions across both sets. This class distribution aligns with real-world SAR imagery characteristics, where certain target types naturally appear more frequently than others due to the inherent nature of SAR remote sensing applications and operational scenarios.
@@
Metric Training Set Test Set Diff-Rate
Mean 1.28 1.26 -0.02
Median 1.062 1.05 -0.017
SD 1.18 0.91 -0.22
@@
As shown in Table 2, the differences in central tendency between training and test sets are minimal (mean: -0.02, median: -0.017). The test set exhibits a 0.22 lower standard deviation, indicating a more concentrated distribution. The key morphological distribution intervals of targets are illustrated in Appendix A.2.
The dataset exhibits three distinct morphological categories based on aspect ratio (AR): broad-bodied (AR $\leq$ 0.67), nearly square-shaped (0.67 $<$ AR $\leq$ 1.5), and tall-bodied (AR $>$ 1.5). Detailed distribution analysis can be found in Appendix A.2.
@@
4 SARChat-Bench Evaluation Method and Settings
This section details the evaluation methodology of SARChat-Bench, a standardized benchmark suite we designed for comprehensive assessment of VLMs in SAR interpretation. The benchmark covers six fundamental tasks that evaluate the model’s core capabilities across information processing, target localization, and semantic understanding, providing multi-dimensional insights into visual-language model performance in the SAR domain. The evaluation framework ensures fair and thorough assessment of VLMs’ capabilities across different SAR interpretation scenarios.
@@ 4.1 Evaluation Metrics
Among them, $n_{t}$ represents the sample size of task $t$, $a_{m,t}$ denotes the accuracy of model $m$ on task $t$, and $T$ is the set of all tasks. The detailed calculation of $a_{m,t}$ for each task can be found in Appendix C.
@@ 4.2 Assessment Methods
(1) Instance Counting: Compare predicted and label object counts for single-class evaluation, where counting accuracy is measured by $Acc$ and object localization precision is evaluated using Intersection over Union ($IoU$).
@@
(6) Classification: Compare predicted and ground-truth categories to assess classification accuracy $Acc$.
@@ 5.1 Implementation Details
During the fine-tuning stage, the model is trained for 1 epoch with a batch size of 4 (effective batch size 32 with gradient accumulation steps of 4). This study employs LoRA training method with rank=8 and alpha=32 on all linear layers. The learning rate is initialized at 1e-4 with a 0.1 warm-up ratio. All experiments are conducted on 2 NVIDIA A100 GPUs using bfloat16 precision. Training is implemented using the MS-SWIFT framework (Zhao et al., 2024) for efficient distributed training. In the most time-consuming case, the training process for a single model took up to 192 hours to complete.
@@ 5.2 Benchmark Evaluation
To verify the effectiveness and practicality of the SARChat-2M dataset, we conducted extensive experiments on SAR image interpretation tasks using sixteen mainstream visual-language models. Our preliminary analysis reveals that these models, despite their strong performance on natural images, struggle significantly with SAR image interpretation without domain-specific fine-tuning, demonstrating the critical importance of SAR-domain adaptation. Specifically, we conducted a detailed before-and-after analysis on InternVL2-8B, which achieved the best performance among all tested models, to quantitatively demonstrate the impact of SAR fine-tuning. The results are presented in Appendix E.
As shown in Table 3, the evaluated models include recent advances such as Qwen2-VL(Wang et al., 2024), InternVL2.5(Chen et al., 2024), DeepSeekVL(Lu et al., 2024), Phi-3.5-vision(Abdin et al., 2024), GLM-Edge-V(GLM et al., 2024), mPLUG-Owl3(Ye et al., 2024), Yi-VL(Young et al., 2024) and LLaVA-1.5(Liu et al., 2023). In experiments of different tasks, each model exhibits diverse characteristics:
@@
Figure 3: Evaluation examples on SARChat-Bench. VLM predictions are shown in green/red for correct/incorrect descriptions, with the ground truth in green and the predictions in red boxes. And [Human], [Bot], and [Check] icons denote user input, VLMs response, and standard output, respectively.
@@ 5.4 Dialogue Visualizaion
Figure 3 presents examples across six tasks from SARChat-Bench. The model completes these tasks with reasonable performance, and its coordinate predictions align with the ground truth annotations. In the spatial grounding task, the model identifies an additional ship that was not included in the original annotations, suggesting its potential in detecting previously unmarked targets in SAR imagery.
@@ 6 Conclusion
This research introduces SARChat-2M, a large-scale dataset of two million annotated SAR image-text pairs, addressing the scarcity of language-vision data in the SAR domain. The accompanying SARChat-Bench provides a systematic evaluation framework for assessing VLMs in SAR interpretation tasks, facilitating domain-specific knowledge integration and accelerating the development of SAR-oriented VLMs.
Limitation
Despite the comprehensive scale of SARChat-2M based on SARDet-100K dataset, the inherent annotation inconsistencies across different SAR data sources may lead to potential limitations. The varying annotation quality could result in missing targets or imprecise target delineation. Notably, there exist cases where VLMs successfully identify valid targets that were not originally annotated in the dataset, highlighting the annotation completeness challenge in the current benchmark construction.
Ethics Statement
In this study, all SAR datasets and methodologies are used strictly for academic research purposes, adhering to their respective licenses and data usage agreements. While our research aims to advance the fundamental understanding of SAR image interpretation, we acknowledge that these technologies could potentially be applied to military or defense-related purposes. We emphasize that the responsible application of such technologies is crucial, and their deployment should strictly comply with relevant regulations and ethical guidelines. The research community should maintain ongoing discussions regarding the dual-use nature of SAR technologies to ensure their development serves beneficial purposes while minimizing potential misuse.
@@
Appendix A Data structure analysis
A.1 Word Frequency Analysis of SARChat-2M
@@
A.2 Analysis of Datasets Composition
@@
Table 4: Analysis of Aspect Ratio of Different Types of Targets
@@
A.3 Analysis of Task based on Datasets Composition
@@
Table 5: Task type distribution in training and test sets
@@ A.4 Dataset Examples
Appendix B Motivation for the Dual-benchmark Setup
@@
Performance 8B 4B 2B 1B
@@
Table 6: Performance comparison across different model sizes
@@
Appendix C Task-specific Performance Scoring
@@
Appendix D The Analysis of Model Size
@@
Model Only count Spatial Ground Cross-Modal ID Multi-target Ref Single-target Ref Descript Class
Multi Single Multi Single IoU=.25 IoU=.5 IoU=.25 IoU=.5
InternVL2.5(8B) 63.71 5.07 0.38 0.12 0.17 0 0 0 0 0.04 17.8
SARChat-InternVL2.5(8B) 74.14 62.25 87.91 98.84 98.98 37.49 23.46 74.86 60.13 63.43 97.25
Table 7: Performance comparison of InternVL2.5 before and after fine-tuning
Figure 10: Simple Examples of ship detection in SAR images. The ships appear as distinct bright spots in these SAR images, making them relatively easy to count even for VLMs without SAR-specific training.
Appendix E Comparison before and after training
To evaluate the impact of the SARChat-2M training dataset, we conducted a comparative analysis on InternVL-2.5-8B—the best-performing model in SARChat-Bench—before and after fine-tuning. Our evaluation metrics reveal that without fine-tuning on SARChat-2M, the model fails to comprehend most SAR-specific targets, exhibiting near-zero performance on tasks involving target interpretation and description. The only exception is the Instance Counting Task, where InternVL2.5-8B achieves a baseline accuracy of 63.71%. This relatively high performance can be attributed to the prevalence of ship-on-sea samples, where SAR imaging exhibits distinctive characteristics: the smooth sea surface creates specular reflection, causing most electromagnetic waves to scatter away from the sensor direction, resulting in weak returns that appear as dark areas. Meanwhile, ships’ metallic structures and geometric features (such as dihedral and trihedral corner reflectors) generate strong backscattering, concentrating radar waves back to the sensor, thus appearing as bright spots. As shown in Figure 10, these samples present relatively straightforward recognition scenarios, leading to higher counting accuracy scores. The comparative results between the base model and its fine-tuned version on SARChat-Bench are presented in Table 7. Overall, fine-tuning with the SARChat-2M dataset proves essential for enabling VLMs to interpret SAR imagery effectively.